# Static Three-Dimensional Structures Determine Fast Dynamics Between Distal Loci Pairs in Interphase Chromosomes

### v2

@@ Abstract
Live cell imaging experiments have shown that, although eukaryotic chromosomes are compact, the distal dynamics between enhancers and promoters are unexpectedly rapid and cannot be explained by standard polymer models. The discordance between the compact static chromatin organization (structure) and the relaxation times (dynamics) is a conundrum that violates the expected structure-function relationship. To resolve the puzzle, we developed a theory to predict chromatin dynamics by first calculating the accurate three-dimensional (3D) structure using static Hi-C contact maps or fixed-cell imaging data. The calculated 3D coordinates are used to accurately forecast the two-point dynamics reported in recent experiments that simultaneously monitor chromatin dynamics and transcription. Strikingly, the theory not only predicts the observed fast enhancer-promoter dynamics but also reveals a novel scaling relationship between two-point relaxation time and genomic separation that is in near quantitative agreement with recent experiments. The theory also predicts that cohesin depletion speeds up the dynamics between distal loci. Both the rapid dynamics between distal loci pairs in interphase chromosomes and the acceleration of chromatin loci diffusion upon cohesin depletion are explained in terms of structure-based centrality measure used in graph theory. Our framework shows that chromatin dynamics can be predicted based solely on static experimental data, reinforcing the concept that the three-dimensional structure determines their dynamic behavior. The generality of the theory opens new avenues for exploring chromatin dynamics, especially transcriptional dynamics, across different biological contexts.
@@ I Introduction
Over the last fifteen years, our understanding of chromatin organization has increased significantly, thanks to advances in experimental techniques, such as Chromosome Conformation Capture and its variants (collectively referred to as Hi-C) [1, 2], as well as multiplexed FISH and other fixed-cell imaging methods [3, 4, 5, 6, 7, 8, 9, 10]. These studies, combined with computational modeling [11, 12, 13, 14, 15, 16, 17, 18, 19], have revealed the organizational principles that underlie three-dimensional structures of chromosomes at both the ensemble (obtained by averaging over a cell population) and single-cell level. For instance, multiplexed-FISH experiments [20, 8, 9] and polymer theory [21] have been used to show that chromosomes exhibit extensive conformational heterogeneity at the single-cell level, reflecting the dynamical nature of their organization. The combination of experiments and polymer modeling has provided insights into the organization of interphase as well as mitotic chromosomes [22, 23].
@@
In a recent significant development, Brückner et al. [36] employed a three-color labeling scheme to simultaneously probe the dynamics of several pairs of enhancers and promoters along with the transcription of the corresponding gene. The key results of the study, which investigated the one-point and two-point dynamical correlations of chromatin in Drosophila cells, may be summarized as follows. (1) On the genomic scale, 58 kb $\leq s\leq$ 3.3 Mb where $s$ is the linear genomic length, chromosomes are compact, resembling the fractal globule (FG) model [37]. This implies that the mean distance, $r(s)$, between two loci should scale as $r(s)\sim s^{\nu}$ where $\nu=1/3$ (fractal dimension is $1/\nu$). (2) However, the diffusion exponent, $\alpha$, characterizing the Mean Square Displacement (MSD) of single chromatin loci and the two-point MSD are both approximately 0.5, which is close to the prediction using the Rouse model [38]. (3) Most notably, the relaxation time ($\tau$), associated with the two-point correlation, scales as a power law, $\tau\sim s^{\gamma}$ where $\gamma\approx 0.7$. Surprisingly, the measured ${\gamma}$ value is significantly smaller than the predictions based on both the Rouse model ($\gamma=2$) and the Fractal Globule (FG) model ($\gamma=5/3$) [37, 1, 39]. It is striking that the relaxation dynamics between pairs of loci occur on time scales that are significantly faster than predictions using dynamic scaling arguments that are based on the estimates of the mean separation between the loci using the FG or the Rouse model. The apparent lack of connection between the global static structures ($r(s)$ as a function of $s$) and the observed dynamic behavior requires a theoretical explanation.
@@
Let us briefly explain the origin of the conundrum noted in the experiments [36]. The typical relaxation time, $\tau_{r}$, of a polymer coil on the scale $s$ (measured along the polymer contour or genomic length) is given by $\tau_{r}=r^{2}(s)/D(s)$, where $r(s)$ is the characteristic size of the polymer at scale $s$, and $D(s)$ is the associated diffusion coefficient. If we assume that $D(s)$ obeys $D(s)\sim s^{-\theta}$ and $r(s)\sim s^{2\nu}$, we obtain the well-known relation $\tau_{r}\sim s^{2\nu+\theta}$ [40]. On the other hand, the time scale for single monomer diffusion at time $\tau_{r}$ must be consistent with $r(s)$, leading to the relation $\tau_{r}^{\alpha}\sim r^{2}(s)\sim s^{2\nu}$. Consequently, the diffusion exponent for a single monomer at intermediate timescales is $\alpha=2\nu/(2\nu+\theta)$. However, $\tau_{r}$ described above is difficult to quantify directly experiments. Instead, Brückner et al. [36] measured the two-point dynamics using $M_{2}(t)=\langle||\bm{r}_{ij}(t)-\bm{r}_{ij}(0)||^{2}\rangle$ and defined the relaxation time $\tau$ as the time at which $M_{2}(\tau)$ saturates at $\langle r^{2}(s)\rangle$. The scaling analysis shows that $\tau$ follows the same dependence on $s$ as $\tau_{r}$, namely, $\tau\sim s^{2\nu+\theta}$. This shows that $\gamma=2\nu+\theta$. The scaling relation,
@@ II Results
Outline of the Theory: We developed a theory based on the supposition that knowledge of the static three-dimensional (3D) structure (namely, the knowledge of all the three-dimensional coordinates, $\{\bm{r}_{i}\}$ of the chromatin loci) is sufficient to accurately predict the dynamics between arbitrary pairs of loci. The theory is executed in two steps. (i) We first use the measured (Hi-C or related methods) contact map to calculate the precise 3D structures [41] based on the maximum entropy principle, which yields the joint distribution function, $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$. The Hi-C contact map is used to calculate the mean distances ($\langle r_{ij}\rangle$) between loci $i$ and $j$ using polymer physics concepts [8, 42]. The values of $\langle r_{ij}\rangle$ are needed to calculate $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$. The Lagrange multipliers (parameters), $k_{ij}$, in Eq. 2, ensure that the mean distances between all pairs of loci match the calculated values using $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$. (ii) By interpreting $k_{ij}$ as spring constants in a harmonic potential in the chromatin network, we calculated the dynamical correlation functions using standard procedures used in the theory of polymer dynamics [38]. The details follow.
3D Structures from Hi-C Data: The first step in the theory is the determination of the ensemble of 3D structures that are quantitatively consistent with the measured contact map. To this end, we used the polymer physics-based HIPPS (Hi-C-Polymer-Physics-Structures) [41] and the related DIMES [43] methods. The HIPPS relates the probability of contact, $\langle p_{ij}\rangle$, between loci $i$ and $j$, and the mean spatial distance $\langle r_{ij}\rangle$ separating them [8, 42] through the power law relation, $\langle r_{ij}\rangle=\Lambda{\langle p_{ij}\rangle}^{-1/\alpha}$ with $\alpha\approx 4$. This relation, which was first noted in imaging experiments [8] and subsequently validated in simulations [42], differs from the predictions based on standard polymer models. With $\langle r_{ij}\rangle$ in hand, we formulate the maximum-entropy distribution as a function of the 3D chromatin loci,
@@
Dynamics from $\bm{K}$: Although the distribution $P^{\mathrm{MaxEnt}}(\bm{r}_{1},\bm{r}_{2},\cdots)$ (Eq. 2) is calculated using the maximum-entropy principle, we interpret it as a Boltzmann distribution at unit temperature ($k_{B}T$ is unity) with an effective energy, $H=\sum_{i<j}k_{ij}||\bm{r}_{i}-\bm{r}_{j}||^{2}$. With this identification, $k_{ij}$ may be interpreted as the spring constant between loci $i$ and $j$. Note that some $k_{ij}$ values are allowed to be negative, which indicates repulsion between chromatin loci. Despite the presence of some negative $k_{ij}$ values, $-\bm{K}$ is positive semidefinite  which ensures that the probability distribution $P^{\mathrm{MaxEnt}}(\bm{r}_{1},\bm{r}_{2},\cdots)$ is well-defined and normalizable. The interpretation that $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$ resembles a Boltzmann distribution allows us to derive the inter-loci dynamics using the framework employed in the context of the Rouse model [38]. Therefore, the eigen-decomposition of the connectivity matrix $\bm{K}$ may be used to calculate the normal modes. Each independent normal mode obeys the Ornstein-Uhlenbeck process. With this assumption, dynamical quantities such as $M_{2}(t)$ can be expressed in terms of the eigenvalues and eigenvectors of $\bm{K}$ (see Supplementary Information for details).
To understand the loci relaxation dynamics, we define the two-point correlation function $G_{2}(t)$ as $G_{2}(t)=\langle r^{2}(s)\rangle-M_{2}(t)/2$. The dynamical scaling form of $G(t)$ should be $G_{2}(t)/G_{2}(0)\sim g(ts^{-b})$. At $t=\tau$, the curves collapse with $\tau s^{-b}\sim 1$ which leads to $b=2\nu+\theta$. It has been shown [44] that for $s\ll N$, the scaling form is, $G_{2}(t)\sim t^{(2\nu-2)/(2\nu+1)}$. $G(t)$ also provides a well-defined way to define the relaxation time $\tau$ using $G(\tau)=1/e$. In our theory, the auto-correlation function for the pair of loci $i$ and $j$, $G_{2}^{ij}(t)$, is given by,
@@
where $p$ is the normal mode index, and $\lambda_{p}$ and matrix $\bm{V}$ are the eigenvalues and eigenvectors of $\bm{K}$, respectively. The structure of Eq. 3 matches the Rouse model dynamics [45, 38], except in the chromatin $\bm{V}$ is non-trivial that requires numerical evaluation using the measured contact maps. The relaxation time for each normal mode is $\tau_{p}=-\xi/\lambda_{p}$ where $\xi$ is the friction coefficient. The two-point MSD, $M_{2}(t)$, is calculated from $G_{2}(t)$ using $M_{2}(t)=2\langle r_{ij}^{2}\rangle-2G_{2}(t)$ where $\langle r_{ij}^{2}\rangle$ is the square of the equilibrium mean spatial distance between the two chromatin loci. Note that evaluation of Eq. 3 requires only the properties of the matrix, $\bm{K}$. In the Rouse model $\bm{K}$ is the polymer connectivity matrix, which is tri-diagonal. In the chromatin problem, it is calculated using Eq. 2 for which the experimental Hi-C or Micro-C contact map is required.
@@
Figure 1: (a) Normalized two-point relaxation time, $G_{2}(t)/G_{2}(0)$, for pair of monomer of indices $i$ and $j$ with $s=|j-i|$. Monomers are selected symmetrically around the center of the chain with $s=|j-i|$ where $i=N/2-s/2$ and $j=N/2+s/2$. The results are for the Rouse model with chain length $N=1,000$. Solid circles indicate the relaxation time $\tau$, defined as $G_{2}(\tau)/G_{2}(0)=1/e$. (b) Relaxation time $\tau$ as a function of the sub-chain length $s$ shows $\tau(s)\sim s^{x}$; $x=2$ for the Rouse model, $x=5/3$ for the Fractal Globule (FG), and $x\approx 2.2$ for self-avoiding walk (SAW) chain. The chain length for all three models is $N=1,000$.
@@
Validating the theory: To validate our theory, let us first show that HIPPS-DIMES (we refer our model as HIPPS-DIMES from now on) correctly recovers the known scaling relations for the Rouse chain, self-avoiding walk (SAW), and FG. The mean spatial distance map for these models can be analytically constructed by using the well-known results, $\langle r_{ij}\rangle=|i-j|^{1/2}$, $\langle r_{ij}\rangle=|i-j|^{3/5}$, and $\langle r_{ij}\rangle=|i-j|^{1/3}$, respectively. Using the analytic expressions for the mean distances, the first step in this theory is to determine $\bm{K}$ so that target mean pairwise distances $\langle r_{ij}\rangle$ are recovered. In the example, we set the total length of chain to be 1,000, and consider two monomers to be symmetrically located around the midpoint separating them by a linear genomic distance, $s$. Using $\bm{K}$, we calculated $G_{2}(t)$ (Eq. 3) for different models. Fig. 1(a) shows the $G_{2}(t)$ for an ideal chain for different $s$ values. The relaxation times $\tau$, obtained using $G_{2}(\tau)/G_{2}(0)=1/e$, are shown as solid circles in Fig. 1(b). Similarly, $G_{2}(t)$ for the SAW and FG are calculated (see Supplementary Figure 3). Fig. 1(b), showing the dependence of $\tau$ as a function of $s$, establishes that the expected scaling of $s^{2}$, $s^{5/3}$, and $s^{2.2}$ are correctly reproduced for an ideal chain, the FG, and the self-avoiding chain [46], respectively. These calculations show that as long as the 3D polymer structures are known then the relaxation times may be accurately calculated. Needless to say that the dependence of $\tau$ with $s$ for homopolymers may be obtained using well-known scaling arguments without resorting to simulations.
@@
To further demonstrate that the theory reproduces the correct dynamical properties, we also tested it against the simulation of self-avoiding polymers in both good and poor solvents. We first performed equilibrium Brownian Dynamics simulations of self-avoiding polymers (see SI for details). We then computed the mean distance matrix from the trajectories. Using the mean distance matrices as input, we calculated the connectivity matrix using the maximum entropy principle (Eq. 2). The connectivity matrix could be used to calculate the single-monomer MSD $M_{1}(t)$ and two-point MSD $M_{2}(t)$. Supplementary Figure 4 and Figure 5 show that the theory accurately reproduces the correct dynamics in both good and poor solvents.
@@
Figure 2: (a) Comparison between the experimental data (Micro-C) and the predictions using the HIPPS-DIMES method for chromosome 2R in Drosophila. The anchor (promoter) and seven different distal loci are represented by squares and circles, respectively. The loci are chosen according to the eve promoter-enhancer experimental setup [36]. (b) Distribution of pairwise distances for the seven promoter-enhancer loci pairs. The bar plots show experimental data from [36], while the solid lines represent HIPPS-DIMES predictions. Jensen-Shannon divergence (JSD) between the empirical distributions and model predictions are shown. (c) Comparison of the mean spatial distances $\langle r\rangle$ as a function of the genomic distance $s$ between the experimental measurements and the theoretical predictions. (d) Two-point Mean Square Displacement ($M_{2}(t)$) calculated from the theory using $M_{2}(t)=2\langle r_{ij}^{2}\rangle-2G_{2}(t)$ with $G_{2}(t)$ from Eq. 3. (e) Comparison of relaxation times $\tau$ as a function of genomic distance $s$ between the experimental observations and theoretical predictions. Black dashed line is our fit to the experimental data. (f). Relaxation time $\tau$ versus the genomic distance $s$ (black line is a fit) for all pairs of loci.
@@
Application to experiments: Having established that the theory correctly reproduces the dynamics of an ideal chain, FG as well as self-avoiding chain, we use it to resolve the conundrum that the equilibrium distances between pairs of loci are incompatible with the observed transcriptional dynamics [36]. The derivation leading to Eq. 3 (see SI for details) shows that, if all the loci in the chromatin experience the same friction coefficient $\xi$, the dynamics based on the HIPPS-DIMES model is fully determined by the connectivity matrix $\bm{K}$. The expression in Eq. 3 can be numerically computed using the eigenvalues/eigenvectors of $\bm{K}$ (details are given in the SI). In the HIPPS-DIMES theory, $\bm{K}$ for any chromosome may be readily calculated from the static contact map (Hi-C or Micro-C) or the imaging data. Since our theory relies on contact map or imaging data to generate the mean distance map, we utilized the published Micro-C data for Drosophila embryo cells [47], the cell line used in the experimental study probing the dynamics [36]. Fig. 2(a) compares the HIPPS-DIMES prediction for the contact map with the Micro-C data. In addition to reproducing the contact map faithfully, Fig. 2(b) shows that the distributions, $P(r)$, of spatial distance between promoter and seven enhancers are quantitatively recovered (Jensen-Shannon Divergence (JSD) between the empirical distributions and model predictions are calculated and shown). Fig. 2(c) shows the spatial distances, $\langle r\rangle$, as a function of $s$. These results show that the predictions of the HIPPS-DIMES, using the Micro-C contact map as input, are in excellent agreement with experiments.

### v3

@@ Abstract
Live-cell imaging experiments have shown that the distal dynamics between enhancers and promoters are unexpectedly rapid and incompatible with standard polymer models. The discordance between the compact static chromatin organization and dynamics is a conundrum that violates the expected structure-function relationship. We developed a theory to predict chromatin dynamics by accurately determining three-dimensional (3D) structures from static Hi-C contact maps or fixed-cell imaging data. Using the calculated 3D coordinates, the theory accurately forecasts experimentally observed two-point chromatin dynamics. It predicts rapid enhancer-promoter interactions and uncovers a scaling relationship between two-point relaxation time and genomic separation, closely matching recent measurements. The theory predicts that cohesin depletion accelerates single-locus diffusion while significantly slowing relaxation dynamics within topologically associating domains (TADs). Our results demonstrate that chromatin dynamics can be reliably inferred from static structural data, reinforcing the notion that 3D chromatin structure governs dynamic behavior. This general framework offers powerful tools for exploring chromatin dynamics across diverse biological contexts.
@@ I Introduction
Over the last fifteen years, our understanding of chromatin organization has increased substantially, thanks to advances in experimental techniques, such as Chromosome Conformation Capture and its variants (collectively referred to as Hi-C) [1, 2], as well as multiplexed FISH and other fixed-cell imaging methods [3, 4, 5, 6, 7, 8, 9, 10]. These studies, combined with computational modeling [11, 12, 13, 14, 15, 16, 17, 18, 19], have revealed the organizational principles that underlie the three-dimensional structures of chromosomes at both the ensemble (obtained by averaging over a cell population) and the single-cell level. For instance, multiplexed-FISH experiments [20, 8, 9] and polymer theory [21] have been used to show that chromosomes exhibit extensive conformational heterogeneity at the single-cell level, reflecting the dynamical nature of their organization. The combination of experiments and polymer modeling has provided insights into the organization of interphase as well as mitotic chromosomes [22, 23].
@@
In a recent notable development, Brückner et al. [36] employed a three-color labeling scheme to simultaneously probe the dynamics of several pairs of enhancers and promoters along with the transcription of the corresponding gene. The key results of the study, which investigated the one-point and two-point dynamical correlations of chromatin in Drosophila cells, may be summarized as follows. (1) On the genomic scale, 58 kb $\leq s\leq$ 3.3 Mb where $s$ is the linear genomic length, chromosomes are compact, resembling the fractal globule (FG) model [37]. This implies that the mean distance, $r(s)$, between two loci should scale as $r(s)\sim s^{\nu}$ where $\nu=1/3$ (fractal dimension is $1/\nu$). (2) However, the diffusion exponent, $\alpha$, characterizing the mean square displacement (MSD) of single chromatin loci and the two-point MSD are both approximately 0.5, which is close to the prediction using the Rouse model [38]. (3) Most notably, the relaxation time ($\tau$), associated with the two-point correlation, scales as a power law, $\tau\sim s^{\gamma}$ where $\gamma\approx 0.7$. Surprisingly, the measured ${\gamma}$ value is substantially smaller than the predictions based on both the Rouse model ($\gamma=2$) and the Fractal Globule (FG) model ($\gamma=5/3$) [37, 1, 39]. It is striking that the relaxation dynamics between pairs of loci occur on time scales that are substantially faster than predictions using dynamic scaling arguments that are based on the estimates of the mean separation between the loci using the FG or the Rouse model. The apparent lack of connection between the global static structures ($r(s)$ as a function of $s$) and the observed dynamic behavior requires a theoretical explanation.
@@
Let us briefly explain the origin of the conundrum noted in the experiment [36]. The typical relaxation time, $\tau_{r}$, of a polymer coil of length $s$ (measured along the polymer contour or genomic length) is given by $\tau_{r}=r^{2}(s)/D(s)$, where $r(s)$ is the characteristic size of the polymer, and $D(s)$ is the associated diffusion coefficient. If we assume that $D(s)$ obeys $D(s)\sim s^{-\theta}$ and $r^{2}(s)\sim s^{2\nu}$, we obtain the well-known relation $\tau_{r}\sim s^{2\nu+\theta}$ [40]. On the other hand, the time scale for single monomer diffusion at time $\tau_{r}$ must be consistent with $r(s)$, leading to the relation $\tau_{r}^{\alpha}\sim r^{2}(s)\sim s^{2\nu}$. Consequently, the diffusion exponent for a single monomer at intermediate timescales is $\alpha=2\nu/(2\nu+\theta)$. However, $\tau_{r}$ described above is difficult to quantify directly through experiments. Instead, Brückner et al. [36] measured the two-point dynamics using $M_{2}(t)=\langle||\bm{r}_{ij}(t)-\bm{r}_{ij}(0)||^{2}\rangle$ where $\bm{r}_{ij}(t)$ is the time-dependent vector pointing from locus $i$ to locus $j$ and defined the relaxation time $\tau$ as the time at which $M_{2}(\tau)$ saturates at $\langle r^{2}(s)\rangle$. The scaling analysis shows that $\tau$ follows the same dependence on $s$ as $\tau_{r}$, namely, $\tau\sim s^{2\nu+\theta}$. This shows that $\gamma=2\nu+\theta$. The scaling relation,
@@ II Results
Outline of the Theory: We developed a theory based on the supposition that knowledge of the static three-dimensional (3D) structure (namely, knowledge of all three-dimensional coordinates, $\{\bm{r}_{i}\}$ of the chromatin loci) is sufficient to accurately predict the dynamics between arbitrary pairs of loci. The theory is executed in two steps. (i) We first use the measured (Hi-C or related methods) contact map to calculate the precise 3D structures [41] based on the maximum entropy principle, which yields the joint distribution function, $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$. The Hi-C contact map is used to calculate the mean distances ($\langle r_{ij}\rangle$) between loci $i$ and $j$ using polymer physics concepts [8, 42]. The values of $\langle r_{ij}\rangle$ are needed to calculate $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$. The Lagrange multipliers (parameters), $k_{ij}$, in Eq. 2, ensure that the mean distances between all pairs of loci match the calculated values using $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$. (ii) By interpreting $k_{ij}$ as spring constants in a harmonic potential in the chromatin network, we calculated the dynamical correlation functions using standard procedures used in the theory of polymer dynamics [38]. The details follow.
3D Structures from Hi-C Data: The first step in the theory is the determination of the ensemble of 3D structures that are quantitatively consistent with the measured contact map. To this end, we used the polymer physics-based HIPPS (Hi-C-Polymer-Physics-Structures) [41] and the related DIMES [43] methods. The HIPPS relates the probability of contact, $\langle p_{ij}\rangle$, between loci $i$ and $j$, and the mean spatial distance $\langle r_{ij}\rangle$ separating them [8, 42] through the power law relation, $\langle r_{ij}\rangle=\Lambda{\langle p_{ij}\rangle}^{-1/\alpha}$ with $\alpha\approx 4$. This relation, which was first reported in imaging experiments [8] and subsequently validated in simulations [42], differs from the predictions based on standard polymer models. To further validate this choice, we compared the inferred mean pairwise distances against imaging data over a range of $\alpha$ values. We find that $\alpha=4$ minimizes the root-mean-square deviation (fig. S11). With DIMES, we directly utilize the imaging data (coordinates of loci) to compute mean pairwise distances. In this work, we will refer our theory as HIPPS-DIMES. With $\langle r_{ij}\rangle$ in hand, we formulate the maximum-entropy distribution as a function of the chromatin loci coordinates,
@@
Dynamics from $\bm{K}$: Although the distribution $P^{\mathrm{MaxEnt}}(\bm{r}_{1},\bm{r}_{2},\cdots)$ (Eq. 2) is calculated using the maximum-entropy principle, we interpret it as a Boltzmann distribution at unit temperature ($k_{B}T$ is unity) with an effective energy, $H=\sum_{i<j}k_{ij}||\bm{r}_{i}-\bm{r}_{j}||^{2}$. With this identification, $k_{ij}$ may be interpreted as the spring constant between loci $i$ and $j$. Note that $k_{ij}$ values are allowed to be negative, which indicates repulsion between chromatin loci. Despite the presence of negative $k_{ij}$ values, by construction $-\bm{K}$ remains positive semidefinite (although excessively large negative perturbations could violate semidefiniteness), thus ensuring that the probability distribution $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$ is well-defined and normalizable. The interpretation that $P^{\mathrm{MaxEnt}}(\{\bm{r}_{i}\})$ resembles a Boltzmann distribution allows us to derive the inter-loci dynamics using the framework employed in the context of the Rouse model [38]. Therefore, the eigen-decomposition of the connectivity matrix $\bm{K}$ may be used to calculate the normal modes. Each independent normal mode obeys the Ornstein-Uhlenbeck process. With this assumption, dynamical quantities such as $M_{2}(t)$ can be expressed in terms of the eigenvalues and eigenvectors of $\bm{K}$ (see Supplementary Materials for details). It should be emphasized that the off-diagonal “spring constants” associated with $\bm{K}$ should only be viewed as effective couplings to enforce the Hi-C or imaging-derived distance constraints. They are not literal molecular forces that act on large length scale.
To understand the loci relaxation dynamics, we define the two-point auto-correlation function $G_{2}(t)$ as $G_{2}(t)=\langle r^{2}(s)\rangle-M_{2}(t)/2$. The dynamical scaling form of $G(t)$ should be $G_{2}(t)/G_{2}(0)\sim g(ts^{-b})$. At $t=\tau$, the curves collapse with $\tau s^{-b}\sim 1$ which leads to $b=2\nu+\theta$. It has been shown [44] that for $s\ll N$, the scaling form is, $G_{2}(t)\sim t^{(2\nu-2)/(2\nu+1)}$. $G(t)$ also provides a well-defined way to define the relaxation time $\tau$ using $G(\tau)=1/e$. In our theory, the auto-correlation function for the pair of loci $i$ and $j$, $G_{2}^{ij}(t)$, is given by,
@@
where $p$ is the normal mode index, and $\lambda_{p}$ and matrix $\bm{V}$ are the eigenvalues and eigenvectors of $\bm{K}$, respectively. The structure of Eq. 3 matches the Rouse model dynamics [45, 38], except in the chromatin $\bm{V}$ is non-trivial that requires numerical evaluation using the measured contact maps. The relaxation time for each normal mode is $\tau_{p}=-\xi/\lambda_{p}$ where $\xi$ is the friction coefficient, where $\xi$ represents an effective friction coefficient that incorporates the medium experienced by chromatin loci in the nucleus. It is the only adjustable parameter in the theory and merely sets the overall time scale in all the dynamical predictions. The two-point MSD, $M_{2}(t)$, is calculated from $G_{2}(t)$ using $M_{2}(t)=2\langle r_{ij}^{2}\rangle-2G_{2}(t)$ where $\langle r_{ij}^{2}\rangle$ is the equilibrium mean squared spatial distance between the two chromatin loci. Note that evaluation of Eq. 3 requires only the properties of the matrix, $\bm{K}$. In the Rouse model, $\bm{K}$ is the polymer connectivity matrix, which is tri-diagonal. In the chromatin problem, it is calculated using Eq. 2 for which the experimental Hi-C/Micro-C contact map or imagined data is required.
@@
Figure 1: Two-point relaxation dynamics in homopolymer models. (a) Normalized two-point auto-correlation function, $G_{2}(t)/G_{2}(0)$, for pair of monomer of indices $i$ and $j$ with $s=|j-i|$. Loci pairs are selected symmetrically around the midpoint of the chain, $i=N/2-s/2$ and $j=N/2+s/2$. The results are for the Rouse model with chain length $N=1,000$. Solid circles mark the relaxation time $\tau$, defined as $G_{2}(\tau)/G_{2}(0)=1/e$. (b) Log-log plot of relaxation time $\tau$ as a function of the sub-chain length $s$ shows $\tau(s)\sim s^{x}$; $x=2$ for the Rouse model, $x=5/3$ for the Fractal Globule (FG), and $x\approx 2.2$ for self-avoiding walk (SAW) chain. The chain length for all three models is $N=1,000$.
@@
Validating the theory: To validate our theory, let us first show that HIPPS-DIMES correctly recovers the known scaling relations for the Rouse chain, self-avoiding walk (SAW), and FG. The mean spatial distance map for these models can be analytically constructed by using the well-known relations, $\langle r_{ij}\rangle=|i-j|^{1/2}$, $\langle r_{ij}\rangle=|i-j|^{3/5}$, and $\langle r_{ij}\rangle=|i-j|^{1/3}$, respectively. Using the analytic expressions for the mean distances, the first step in this theory is to determine $\bm{K}$ so that target mean pairwise distances $\langle r_{ij}\rangle$ are recovered. The $\bm{K}$ matrices were obtained using an iterative optimization algorithm described in a previous work [43]. The $\bm{K}$ matrices plotted in fig. S3 show the characteristic banded structure for each polymer model, and we confirmed that each inferred $\bm{K}$ exactly enforces the input Flory scaling exponent $\nu$ (=1/2, 3/5, or 1/3) (fig. S3). In this example, we set the total length of chain to be 1,000, and consider two monomers to be symmetrically located around the midpoint separating them by a linear genomic distance, $s$. Using $\bm{K}$, we calculated $G_{2}(t)$ (Eq. 3) for different models. Fig. 1(a) shows the $G_{2}(t)$ for an Rouse chain for different $s$ values. The relaxation times $\tau$, obtained using $G_{2}(\tau)/G_{2}(0)=1/e$, are shown as solid circles in Fig. 1(b). Similarly, $G_{2}(t)$ for the SAW and FG are calculated (see fig. S4). Because we assume a purely diagonal mobility tensor with no hydrodynamic (Zimm) coupling, our model fixes $\theta=1$. Hence, the scaling relation $\tau_{p}\sim s^{2\nu+\theta}$ becomes $\tau\sim s^{2\nu+1}$, thus recovering $\tau\sim s^{2}$ for $\nu=1/2$ (Rouse chain) and $\tau\sim s^{5/3}$ for $\nu=1/3$ (FG). The full Zimm mobility would give $\theta=\nu$ and thus $\gamma=3\nu$. Fig. 1(b), showing the dependence of $\tau$ as a function of $s$, establishes that the expected scaling of $s^{2}$, $s^{5/3}$, and $s^{2.2}$ are correctly reproduced for a Rouse chain, the FG, and the self-avoiding chain [46], respectively. These calculations show that as long as the 3D polymer structures are known then the relaxation times may be accurately calculated. Needless to say that the dependence of $\tau$ with $s$ for homopolymers may be obtained using well-known scaling arguments without resorting to simulations.
@@
To further demonstrate that the theory reproduces the correct dynamical properties, we also tested it against the polymer simulations of self-avoiding polymers in both good and poor solvents. We first performed equilibrium Brownian Dynamics simulations of self-avoiding polymers (see Supplementary Materials for details). We then computed the mean distance matrix from the trajectories. Using the mean distance matrices as input, we calculated the connectivity matrix using the maximum entropy principle (Eq. 2). The connectivity matrix could be used to calculate the single-monomer MSD $M_{1}(t)$ and two-point MSD $M_{2}(t)$. figs. S5 and S6 show that the theory accurately reproduces the correct dynamics in both good and poor solvents.
[…]
@@
figure HIPPS-DIMES accurately captures the static and dynamic properties of Drosophila enhancer–promoter pairs. (a) Comparison between the Micro-C contact map and the predictions using the HIPPS-DIMES method for chromosome 2R in Drosophila, marking the promoter (square) and seven different distal enhancers (circles), chosen according to the eve promoter-enhancer experimental setup [36]. (b) Distribution of pairwise distances for the seven promoter-enhancer loci pairs: experimental histograms (bars: data from Ref. [36]) and HIPPS-DIMES predicted distribution (solid lines). Jensen-Shannon divergence (JSD) values between model predictions and experimental data are reported. (c) Comparison of the mean spatial distances $\langle r\rangle$ as a function of the genomic distance $s$ between the experimental measurements and the HIPPS-DIMES predictions. (d) Two-point Mean Square Displacement ($M_{2}(t)$), calculated using $M_{2}(t)=2\langle r_{ij}^{2}\rangle-2G_{2}(t)$ with $G_{2}(t)$ from Eq. 3. (e) Comparison of relaxation times $\tau$ as a function of genomic distance $s$ between the experimental observations and HIPPS-DIMES predictions. Black dashed line (Orange dashed line) is the fit to the experimental data (HIPPS-DIMES prediction). (f) Scatter plot for the relaxation time $\tau$ versus genomic separation $s$ for all pairs of loci. Power-law fit is shown in black line.
@@
Application to experiments: Having established that the theory correctly reproduces the dynamics of a Rouse chain, FG as well as self-avoiding chain, we use it to resolve the conundrum that the equilibrium distances between pairs of loci are incompatible with the observed transcriptional dynamics [36]. In the Brückner et al. system [36], seven engineered enhancer–promoter (E–P) pairs were assayed in Drosophila embryos at nuclear cycle 14 (nc14): a minimal eve promoter (MS2-tagged) at a fixed locus and a synthetic eve enhancer (ParS-tagged) inserted at defined distances (58 kb, 82 kb, 149 kb, etc.) from the promoter. To mimic each reporter pair, we use the wild-type Micro-C contact map for Drosophila embryo nc14 cells [47] and select the locus at the same genomic positions from the endogenous eve promoter as in the experiment, assuming that a single small insertion does not noticeably alter large-scale contacts. The derivation leading to Eq. 3 (see Supplementary Materials for details) shows that, if all the loci in the chromatin experience the same friction coefficient $\xi$, the dynamics based on the HIPPS-DIMES model is fully determined by the connectivity matrix $\bm{K}$. The expression in Eq. 3 can be numerically computed using the eigenvalues/eigenvectors of $\bm{K}$ (details are given in the Supplementary Materials). In the HIPPS-DIMES theory, $\bm{K}$ for any chromosome may be readily calculated from the static contact map (Hi-C or Micro-C) or the imaging data. Fig. II(a) compares the HIPPS-DIMES prediction for the contact map with the Micro-C data (see fig. S12 for zoom-in view on one Mb-scale comparison). The enhancers and promoter used in the experiment setup [36] are marked. In addition to reproducing the contact map faithfully, Fig. II(b) shows that the distributions, $P(r)$, of spatial distance between promoter and seven enhancers are quantitatively recovered (Jensen-Shannon Divergence (JSD) between the empirical distributions and model predictions are calculated and shown). Model predictions in Fig. II(b) were generated by sampling 10,000 independent HIPPS-DIMES conformations and computing all pairwise distances. Note that the 82 kb (orange) and 88 kb (green) traces overlap exactly due to the 12.812.812.8\,12.8kb genomic-resolution of the reconstructed structures. Fig. II(c) shows the spatial distances, $\langle r\rangle$, as a function of $s$. These results show that the structural predictions of the HIPPS-DIMES, using the Micro-C contact map as input, are in excellent agreement with both Micro-C and imaging experiments.